OpenStudy (anonymous):

Find two consecutive odd numbers such that the sum of the first number and one-third of the second number is equal to fifty. 35&37 37&39 49&51

OpenStudy (zzr0ck3r):

consecutive odd numbers \(n, n+2\) where \(n\) is the first odd number in question. \(n+\frac{1}{3}(n+2)=50\) can you solve for n

OpenStudy (anonymous):

b ?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (triciaal):

yes

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

n=37 n+2 = 39

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

npz